### v2

# Constrained Design of a Binary Instrument in a Partially Linear Model

Abstract
We study the question of how best to assign an encouragement in a randomized encouragement study. In our setting, units arrive with covariates, receive a nudge toward treatment or control, acquire one of those statuses in a way that need not align with the nudge, and finally have a response observed. The nudge can be seen as a binary instrument that affects the response only via the treatment status. Our goal is to assign the nudge as a function of covariates in a way that best estimates the local average treatment effect (LATE). We assume a partially linear model, wherein the baseline model is non-parametric and the treatment term is linear in the covariates. Under this model, we outline a two-stage procedure to consistently estimate the LATE. Though the variance of the LATE is intractable, we derive a finite sample approximation and thus a design criterion to minimize. This criterion is convex, allowing for constraints that might arise for budgetary or ethical reasons. We prove conditions under which our solution asymptotically recovers the lowest true variance among all possible nudge propensities. We apply our method to a semi-synthetic example involving triage in an emergency department and find significant gains relative to a regression discontinuity design.
1 Introduction
Though unconfoundedness is a standard assumption in the causal inference literature, it is easily violated in a great variety of settings. For instance, a scholarship committee may award their finite supply of scholarships to the students who they perceive will derive the most benefit via some impressions not captured in the covariate data. This creates a dependence between treatment status and potential outcomes even conditional on observed covariates that renders many standard causal methods inaccurate.
Consequently, a common practice in the presence of confounding is to run a randomized encouragement design (West et al., 2008). Such studies introduce a binary encouragement, by which some subjects are randomly encouraged to get the treatment (e.g., to apply for the scholarship) and others are not. As long as the encouragement only affects outcomes via its effect on treatment status, this procedure enables certain causal estimands to be identified even under confounding. In particular, a binary encouragement is a kind of binary instrumental variable (IV), enabling IV methods to be used for analysis.
In this paper, we study the effect of a binary encouragement to treatment from a design perspective. That is, supposing that the experimenter gets to choose the encouragement status for each subject – or, more generally, their encouragement probability as a function of their covariates – we consider the question of how they should do so to best learn a treatment effect of interest. Our method also allows for the presence of constraints on randomization, which may arise for budgetary, ethical, or economic reasons. For example, a university may have only finitely many scholarships to give out, and an experiment run in a medical context may need to satisfy certain fairness constraints on whose encouragement probabilities are high or low.
We are motivated by a problem of admission triage for emergency department (ED) patients. There, a decision must be made about which patients should go to intensive care units (ICUs). Some patients clearly must go and others should not. Randomization around any given cutoff would give more information later about which patients benefit most. It is imperative to leave the final decision about ICU admission to the ED physicians along with the patient and family members, and so the treatment itself cannot be randomized. Therefore, our motivating setting involves making a recommendation about which patients should go to ICUs, with the understanding that this recommendation will not always be followed. The recommendation we consider is based on a machine learning-derived prediction of whether a newly admitted patient is likely to go to the ICU. When that estimated probability is very high, it is natural that we might recommend admission to ICUs more often. There are several outcomes of interest in this setting, with the length of hospital stay being prominent among them. More details about this motivating example are given in Section 4.
We study efficiency using a partially linear model, with a non-parametric baseline mean response and a linear model for the treatment effect. This allows us to obtain a finite-sample design criterion that can be optimized under constraints. We do sacrifice some flexibility relative to fully non-parametric IV work based on deep learning, such as Hartford et al. (2017); Chernozhukov et al. (2017); Bennett et al. (2019); Syrgkanis et al. (2019). However, our view is that the partially linear model retains desirable flexibility in the baseline model while giving a concrete and practical criterion for the design stage.
By adapting some techniques for the partially linear model from Robinson (1988), we derive a method of estimating the local average treatment effect (LATE) in our setting. Crucially, the variance of our estimator can be approximated by a quantity that is convex in the encouragement probability, which is what the experimenter is allowed to choose. This enables the construction of a convex optimization procedure and a notion of the optimal encouragement under convex constraints. We show as well that, under certain assumptions presented in Section 3, these approximations to the true variance become negligible as the sample size tends to infinity.
The problem of optimally choosing a propensity score to estimate a treatment effect of interest has been studied in a variety of settings. These include under a linear model with design constraints (Li and Owen, 2023a; Morrison and Owen, 2024), for batch-adaptive designs (Hahn et al., 2011; Li and Owen, 2023b), and for sequential designs (van der Laan, 2008; Kato et al., 2021). However, this question remains minimally explored in the context of confounding and instrumental variables. Several papers, e.g., Kuersteiner and Okui (2010); Kuang et al. (2020), consider the problem of optimally combining a set list of candidate instruments, which is conceptually similar to our question of how best to design a single instrument.
Most similar is the recent paper of Chandak et al. (2023), who study how to choose an optimal policy for instrumental variable assignment in some policy class. Their approach combines the Deep IV framework of Hartford et al. (2017) with the use of influence functions to obtain an estimate of the gradient of the loss with respect to the policy parameters. While addressing the same problem, their method does not yield a convex objective and so has no guarantee of finding the global optimum. In addition, they do not consider the case in which there are constraints on the design, which is important for our motivating example.
This paper is organized as follows. Section 2 introduces the key notation, assumptions, and parameters of interest. Section 3 derives our proposed method and states some relevant theoretical guarantees. Section 4 applies our method to a semi-synthetic example based on an emergency department dataset. Some proofs are in Appendix A. Appendix B generalizes our convexity result to a setting with bounded heteroscedasticity. Appendix C gives details of our simulation.
2 Setup
We use $X_{i}\in\mathcal{X}\subseteq\mathbb{R}^{d}$ to denote the covariates of unit $i$ and $W_{i}\in\{0,1\}$ to denote their (actual) treatment status. Here, $W_{i}=0$ and $W_{i}=1$ correspond respectively to control and treatment. We let $Y_{i}\in\mathbb{R}$ denote the response of interest for unit $i$.
We adopt the potential outcomes framework (Rubin, 1974). We also make the stable unit treatment value assumption (SUTVA) that a subject’s response depends on the vector $W$ only through their own $W_{i}$, i.e., $Y_{i}(W)=Y_{i}(W_{i})$.
Importantly, however, we do not make the standard unconfoundedness assumption that a unit’s treatment status is conditionally independent of its response given $X$. Concretely, if we write
$Y=g(X,W)+\varepsilon,$ (1)
where $\mathbb{E}[\varepsilon]=0$, then as in Hartford et al. (2017), we allow for $\mathbb{E}[\varepsilon\text{ }|\text{ }X,W]\neq 0$.
Instead, we assume access to a binary instrumental variable $Z\in\{0,1\}$, which only impacts $Y$ through its effect on the treatment status $W$. $Z$ can thus be thought of as a binary “nudge” toward receiving treatment. Such instruments arise by construction in randomized encouragement designs, in which the experimenter can only encourage a subject to have a certain treatment status but cannot choose it directly (Bradlow, 1988; Imai et al., 2013). We write $W_{i}(Z_{i})$ for $Z_{i}\in\{0,1\}$ to denote the treatment status that subject $i$ would have if given the nudge status $Z_{i}$. We assume that the nudges $Z_{i}$ are independent but not identically distributed Bernoulli random variables.
We are now ready to state formally the conditions on our instrument $Z$.
Definition 1.
An instrumental variable $Z\in\{0,1\}$ is any random variable that satisfies:
  1. 1.
    Exclusion restriction: $Y(w,z)=Y(w)$ for all $(w,z)\in\{0,1\}^{2}$.
  2. 2.
    Conditional exogeneity: $Z\perp\!\!\!\perp\{Y(0),Y(1),W(0),W(1)\}\text{ }|\text{ }X$.
  3. 3.
[…]
Of note, our assumptions on the instrument $Z$ are conditional on $X$, as in Abadie (2003). This is because, in our motivating setting, we allow for the distribution of $Z$ to depend on $X$.
The first condition states that the outcome $Y$ only depends on the nudge through the nudge’s effect on the treatment. While often hard to justify in observational settings, the example of a nudge in an encouragement design is more compelling, since a nudge is unlikely to impact the outcome via other plausible mechanisms. The second condition states that, conditional on $X$, the nudge is assigned in a way that is independent of the potential outcomes for $W$ and $Y$. In this paper, $Z$ is sampled as a Bernoulli random variable whose probability depends only on $X$, so this assumption holds for free. Finally, the third assumption states that, at any level of $X$, the nudge has some nonzero impact on treatment probability, which is necessary for subsequent identifiability statements.
Because the instrument is binary, our analysis will hinge on a subject’s compliance class, as in Angrist et al. (1996); Abadie (2003). We define a subject to be a complier (C) if $W(0)=0$ and $W(1)=1$, an always-taker (AT) if $W(0)=W(1)=1$, a never-taker (NT) if $W(0)=W(1)=0$, and a defier (D) if $W(0)=1$ and $W(1)=0$. The latter case is implausible in many settings, since it corresponds to a unit actively doing the opposite of whatever they are encouraged to do. For this reason, and because defiers would render it impossible to estimate rates of each compliance status, we will assume that there are no defiers. This assumption is often known as monotonicity and is standard in the IV literature (Imbens and Angrist, 1994). Note that in some settings, treatment is only open to those receiving the nudge and then there are no defiers. The present setting is one where we need to assume that there are no defiers.
For our motivating example, a subject corresponds to an ED patient admitted to the hospital. An always-taker would then be a case severe enough that the physician sends them to an ICU regardless of their nudge status. Analogously, a never-taker is likely a mild case, and a complier is likely an intermediate case.
The compliance status of an observation depends on $X$ but, in our model, is also random. We define the compliance probabilities $p_{\mathrm{C}}(X)=\mathbb{P}(\text{complier}\text{ }|\text{ }X)$, with $p_{\mathrm{AT}}(X)$, $p_{\mathrm{NT}}(X)$ and $p_{\mathrm{D}}(X)$ defined analogously for always-takers, never-takers and defiers, respectively. The relevance criterion in Definition 1 can only be satisfied if $p_{\mathrm{C}}(X)+p_{\mathrm{D}}(X)>0$ at all $X$.
Assumption 1.
(No defiers) For all $X$, $p_{\mathrm{D}}(X)=0$.
We will need to assume that there are some compliers at every $X$. While that follows from Assumption 1 and the relevance condition in Definition 1, we list it as an assumption in its own right.
Assumption 2.
(Some compliers) For all $X$, $p_{\mathrm{C}}(X)>0$.
When there is noncompliance, then the average treatment effect (ATE), $\tau=\mathbb{E}[Y(1)-Y(0)]$, is no longer identified. Instead, a standard target of interest (Imbens and Angrist, 1994; Angrist et al., 1996) is the local average treatment effect (LATE):
$\tau_{\mathrm{LATE}}=\mathbb{E}[Y(1)-Y(0)\text{ }|\text{ }\text{complier}].$
The target $\tau_{\mathrm{LATE}}$ corresponds to the average effect of treatment on those subjects who do respond to the nudge $Z$. Under our assumptions on $Z$, $\tau_{\mathrm{LATE}}$ is identified via the formula
$\tau_{\mathrm{LATE}}=\mathbb{E}_{X}\left[\frac{\mathbb{E}[Y\text{}|\text{ }X,Z=1]-\mathbb{E}[Y\text{ }|\text{ }X,Z=0]}{\mathbb{E}[W\text{ }|\text{ }X,Z=1]-\mathbb{E}[W\text{ }|\text{ }X,Z=0]}\right].$ (2)
This result is a consequence of Theorem 3.1 in Abadie (2003), though we reprove it in Appendix A.1, for completeness.
Figure 1: Illustration of the directed acyclic graph for our problem. The treatment $W$ may be affected by unobserved confounders $U$ (blue arrow), but there exists a binary instrument $Z$ that affects the outcome $Y$ only via its effect on $W$. The dashed red arrows indicate forbidden relationships involving $Z$. Figure inspired by a similar diagram from Chandak et al. (2023).
Of note, the denominator of 2 is precisely $p_{\mathrm{C}}(X)$, since the first expectation in the denominator accounts for both compliers and always-takers whereas the second term accounts for only always-takers.
In this paper, we study the effect of the nudge distribution $Z\text{ }|\text{ }X$ on efficient estimation of $\tau_{\mathrm{LATE}}$. We consider a scenario in which subjects arrive with known covariates $X_{i}\in\mathcal{X}\subseteq\mathbb{R}^{d}$ and unknown compliance status $C_{i}\in\{\text{AT},\text{NT},\text{C}\}$ drawn IID from some distribution $P_{X\times C}$. Subjects are given a nudge $Z_{i}\in\{0,1\}$, then receive a treatment $W_{i}\in\{0,1\}$, and finally have a response $Y_{i}\in\mathbb{R}$. The treatment may be confounded with the unobserved noise that also affects the response, as in Figure 1.
The only aspect of this pipeline that the experimenter can vary is the nudge assignment probability $e_{Z}(X)=\mathbb{P}(Z=1\text{ }|\text{ }X)$. The true propensity score $e_{W}(X)=\mathbb{P}(W=1\text{ }|\text{ }X)$ is not chosen by the experimenter due to noncompliance, but it is affected by $e_{Z}(X)$, since
$e_{W}(X)$ $=\mathbb{P}(W=1\text{ }|\text{ }X)$
$=\mathbb{P}(\text{always-taker}\text{ }|\text{ }X)+\mathbb{P}(\text{complier, }Z=1\text{ }|\text{ }X)$
$=p_{\mathrm{AT}}(X)+p_{\mathrm{C}}(X)e_{Z}(X).$ (3)
3 Methodology
In this section, we define our model with a linear treatment effect on top of a non-parametric baseline. That makes the LATE linear in the parameter vector, allowing us to use methods from optimal design of experiments. We adapt a modeling approach from Robinson (1988) to our IV setting and formulate a constrained optimization problem for the nudges. We give conditions under which the variance of the LATE using a pilot study converges to the true variance, justifying optimal design with pilot estimates. We present an algorithm and note some properties of the resulting optimal designs.
3.1 Partially linear model
To arrive at a criterion for optimality of $e_{Z}(X)$, we will use a more specific model than Equation (1). We choose a partially linear model:
$Y=f(X)+WX^{\top}\gamma+\varepsilon.$ (4)
Here, $f$ is a non-parametric baseline model, $\gamma\in\mathbb{R}^{d}$ is a vector of coefficients for the treatment term, and as in Equation (1), we allow for $\mathbb{E}[\varepsilon\text{ }|\text{ }X,W]\neq 0$ due to the possibility of confoundedness.
This is a well-studied model in the causal inference literature; see Robinson (1988) and Wager (2020). A variant of this model has also been studied in the instrumental variables (IV) setting in, for example, Florens et al. (2012); Chen (2021) but not, to our knowledge, from a design perspective. For our purposes, it has the desirable property of allowing for flexibility in the baseline model while still preserving some parametric structure in the treatment effect that lends itself to a design criterion.
Conveniently, the conditional LATE in the partially linear model (4) is $X^{\top}\gamma$, whereas this is not the case for the conditional ATE due to confoundedness. To see this, observe that
$\tau_{\mathrm{LATE}}$ $=\mathbb{E}_{X}\left[\frac{\mathbb{E}[Y\text{ }|\text{ }X,Z=1]-\mathbb{E}[Y\text{ }|\text{ }X,Z=0]}{p_{\mathrm{C}}(X)}\right]$
$=\mathbb{E}_{X}\left[\frac{(\mathbb{E}[W\text{ }|\text{ }X,Z=1]-\mathbb{E}[W\text{ }|\text{ }X,Z=0])X^{\top}\gamma}{p_{\mathrm{C}}(X)}\right]$
$\quad+\mathbb{E}\left[\frac{\mathbb{E}[\varepsilon\text{ }|\text{}X,Z=1]-\mathbb{E}[\varepsilon\text{ }|\text{ }X,Z=0]}{p_{\mathrm{C}}(X)}\right]$
$=\mathbb{E}_{X}\left[\frac{p_{\mathrm{C}}(X)X^{\top}\gamma}{p_{\mathrm{C}}(X)}\right]$
$=\mathbb{E}[X^{\top}\gamma].$
The crucial property is that $\mathbb{E}[\varepsilon\text{ }|\text{ }X,Z]=\mathbb{E}[\varepsilon\text{ }|\text{ }X]$ because of the conditional exogeneity of the instrument. If we have some estimator $\hat{\gamma}$ for $\gamma$, this immediately suggests $\bar{X}_{n}^{\top}\hat{\gamma}$ as an estimator of $\tau_{\mathrm{LATE}}$, where $\bar{X}_{n}\in\mathbb{R}^{d}$ is the vector of column means of the data matrix $X\in\mathbb{R}^{n\times d}$. Moreover, a natural variance criterion to minimize is then
$\text{Var}(\bar{X}_{n}^{\top}\hat{\gamma}\text{ }|\text{ }X)=\bar{X}_{n}^{\top}\text{Var}(\hat{\gamma}\text{ }|\text{ }X)\bar{X}_{n}.$ (5)
We condition on $X$ in (5) because the $X$ data is not chosen in our setting but rather given. A design that minimizes 5 satisfies a particular case of $C$-optimality (Atkinson et al., 2007), which minimizes the quadratic form of a vector $c$ and a covariance matrix. Here, $c=\bar{X}_{n}$. Our machinery works for other design criteria based on $\text{Var}(\hat{\gamma}\text{ }|\text{ }X)$, such as A-optimality or D-optimality. Note that while $\bar{X}_{n}$ averages over all $n$ subjects regardless of their compliance status, $\bar{X}_{n}^{\top}\hat{\gamma}$ estimates the average treatment effect on just the compliers, i.e., the LATE.
3.2 Estimating the LATE
We now outline a strategy to estimate $\gamma$ in the partially linear model 4. We will assume access to a pilot study of size $n_{0}$ that has been run with some nudge policy that we denote $e_{Z}^{0}(X)$. We then have access to $\{X_{i},Z_{i},W_{i},Y_{i}\}_{i\leq n_{0}}$. This population is necessary to estimate certain nuisance terms, without which we would be unable to develop a design criterion for a nudge policy $e_{Z}$ to estimate $\gamma$ most efficiently. It is analogous to the use of two-stage estimators common in the IV literature (Angrist and Imbens, 1995; Hartford et al., 2017). In practice, $e_{Z}^{0}(X)$ would likely have to satisfy any design constraints imposed on the optimal solution as well. It may also be better to update $e_{Z}(X)$ sequentially or in an online setting rather than assuming two batches of data, but we opt for the latter approach because the methodological differences are minimal. Furthermore, frequent changes to the policy would be disruptive in many settings.
From the pilot data $\{X_{i},Z_{i},W_{i},Y_{i}\}_{i\leq n_{0}}$, we estimate a few nuisance functions. First, we estimate the compliance probabilities $\{p_{\mathrm{C}}(X),p_{\mathrm{AT}}(X),p_{\mathrm{NT}}(X)\}$. At first glance, this may seem difficult because there are some subjects whose compliance status is unknown; a unit with $(Z,W)=(1,1)$, for instance, could be either a complier or an always-taker. However, the assumption of no defiers guarantees that, for example, a unit with $(Z,W)=(0,1)$ is an always-taker. We can then proceed by fitting a categorical regression to estimate $\mathbb{P}_{0}((Z,W)\text{ }|\text{ }X)$ for all $(Z,W)\in\{0,1\}^{2}$, where the subscript 00 emphasizes that these probabilities depend on the particular pilot scheme $e_{Z}^{0}(X)$. Note that
$\mathbb{P}_{0}((Z,W)=(0,1)\text{ }|\text{ }X)$ $=\mathbb{P}_{0}(Z=0\text{ }|\text{ }X)p_{\mathrm{AT}}(X),$
$\mathbb{P}_{0}((Z,W)=(1,0)\text{ }|\text{ }X)$ $=\mathbb{P}_{0}(Z=1\text{ }|\text{ }X)p_{\mathrm{NT}}(X).$
We thus take
$\hat{p}_{\mathrm{AT}}(X)$ $=\frac{1}{1-e_{Z}^{0}(X)}\hat{\mathbb{P}}_{0}((Z,W)=(0,1)\text{ }|\text{ }X),$
$\hat{p}_{\mathrm{NT}}(X)$ $=\frac{1}{e_{Z}^{0}(X)}\hat{\mathbb{P}}_{0}((Z,W)=(1,0)\text{ }|\text{ }X),\quad\text{and}$
$\hat{p}_{\mathrm{C}}(X)$ $=1-\hat{p}_{\mathrm{AT}}(X)-\hat{p}_{\mathrm{NT}}(X).$
In addition, we use the pilot study to estimate the conditional mean functions
$m(X,Z,W)$ $=\mathbb{E}[Y\text{ }|\text{ }X,Z,W]=f(X)+WX^{\top}\gamma+\mathbb{E}[\varepsilon\text{ }|\text{ }X,Z,W]$
$m(X,Z)$ $=\mathbb{E}[Y\text{ }|\text{ }X,Z]=f(X)+(p_{\mathrm{AT}}(X)+p_{\mathrm{C}}(X)Z)X^{\top}\gamma+\mathbb{E}[\varepsilon\text{ }|\text{ }X].$
To avoid ambiguity, the arguments to $m$ will always be explicitly given.
Now suppose that we are considering using the candidate scheme $e_{Z}(X)$ on the main data set. Recall from equation (2) that our choice of a nudge propensity function $e_{Z}$ induces the treatment propensity $e_{W}(X)=p_{\mathrm{AT}}(X)+p_{\mathrm{C}}(X)e_{Z}(X)$. Next, define
$\begin{split}m^{*}(X,Z,W)&=m(X,Z,W)+\frac{e_{W}(X)-W}{p_{\mathrm{C}}(X)}(m(X,Z=1)-m(X,Z=0))\\ &=f(X)+e_{W}(X)X^{\top}\gamma+\mathbb{E}[\varepsilon\text{ }|\text{ }X,Z,W].\end{split}$ (6)
Then
$Y-m^{*}(X,Z,W)=(W-e_{W}(X))X^{\top}\gamma+\varepsilon-\mathbb{E}[\varepsilon\text{ }|\text{ }X,Z,W].$ (7)
If we knew the oracle $m^{*}(X,Z,W)$ and $e_{W}(X)$, then (7) implies that we could recover an unbiased and consistent estimate of $\gamma$ via OLS:
$\hat{\gamma}_{\mathrm{oracle}}=\bigl{(}X^{\top}D(X,W)^{2}X\bigr{)}^{-1}X^{\top}D(X,W)\bigl{(}Y-m^{*}(X,Z,W)\bigr{)}$ (8)
where $D(X,W)=\mathrm{diag}(W-e_{W}(X))$. In practice, we replace these oracle functions with our estimates based on the pilot study:
$\hat{\gamma}_{\mathrm{approx}}=\bigl{(}X^{\top}\hat{D}(X,W)^{2}X\bigr{)}^{-1}X^{\top}\hat{D}(X,W)\bigl{(}Y-\hat{m}^{*}(X,Z,W)\bigr{)},$ (9)
where $\hat{D}(X,W)=\mathrm{diag}(W-\hat{e}_{W}(X))$ for $\hat{e}_{W}(X)=\hat{p}_{\mathrm{AT}}(X)+\hat{p}_{\mathrm{C}}(X)e_{Z}(X)$ and
$\hat{m}^{*}(X,Z,W)=\hat{m}(X,Z,W)+\frac{\hat{e}_{W}(X)-W}{\hat{p}_{\mathrm{C}}(X)}(\hat{m}(X,Z=1)-\hat{m}(X,Z=0).$ (10)
This approach is analogous to ideas in Robinson (1988) but adapted to our IV setting. In the simpler setting without confounding, $Y-\mathbb{E}[Y\text{ }|\text{ }X=x]$ can be regressed on $(W-e_{W}(X))X$ to recover an estimate of the vector $\gamma$. Here, it is necessary instead to subtract off the more complicated $m^{*}(X,Z,W)$ function defined in 6, since we need the error term to be mean zero given $X$, $Z$, and $W$.
3.3 Designing for nudge assignment
To progress toward a design criterion for choosing $e_{Z}(X)$, assume for now that $\text{Var}(\varepsilon-\mathbb{E}[\varepsilon\text{ }|\text{ }X,Z,W]\text{ }|\text{ }X,Z,W)=\sigma^{2}$ is constant in $X$, $Z$, and $W$. We discuss generalizations to heteroscedasticity at the end of this section and in further detail in Appendix B. Then
$\text{Var}(\hat{\gamma}_{\mathrm{oracle}}\text{ }|\text{ }X,Z,W)=\sigma^{2}(X^{\top}\mathrm{diag}(W-e_{W}(X))^{2}X)^{-1}.$
We thus define the oracle variance, which is conditional on only $X$, as
$V(e_{Z})$ $=\text{Var}(\sqrt{n}\,\bar{X}_{n}^{\top}\hat{\gamma}_{\mathrm{oracle}}\text{ }|\text{ }X)$
$=n\bar{X}_{n}^{\top}\mathbb{E}_{Z,W}\bigl{[}\text{Var}(\hat{\gamma}_{\mathrm{oracle}}\text{ }|\text{ }X,Z,W)\bigr{]}\bar{X}_{n}+n\bar{X}_{n}^{\top}\text{Var}(\mathbb{E}[\hat{\gamma}_{\mathrm{oracle}}\text{ }|\text{ }X,Z,W])\bar{X}_{n}$
$=n\sigma^{2}\bar{X}_{n}^{\top}\mathbb{E}_{Z,W}\bigl{[}(X^{\top}\mathrm{diag}(W-e_{W}(X))^{2}X)^{-1}\bigr{]}\bar{X}_{n}.$
Here, we use that $\hat{\gamma}_{\mathrm{oracle}}$ is unbiased for $\gamma$ to eliminate the second term in the law of total variance. We also scale by $\sqrt{n}$ since the unscaled variance in OLS converges to zero. Because $\sigma^{2}$ is merely a multiplicative constant, we take $\sigma^{2}=1$ going forward without loss of generality.
With the above assumptions, a technical difficulty remains because there is a matrix inverse inside of an expectation. To handle it, we make our first of two approximations by moving the inverse out of the expectation:
$V(e_{Z})$ $=n\bar{X}_{n}^{\top}\mathbb{E}_{Z,W}[(X^{\top}\mathrm{diag}(W-e_{W}(X))^{2}X)^{-1}]\bar{X}_{n}$
$\approx n\bar{X}_{n}^{\top}\mathbb{E}_{Z,W}[X^{\top}\mathrm{diag}(W-e_{W}(X))^{2}X]^{-1}\bar{X}_{n}$
$=n\bar{X}_{n}^{\top}(X^{\top}\mathrm{diag}(e_{W}(X)(1-e_{W}(X))X)^{-1}\bar{X}_{n}$
$=V_{\mathrm{approx}}(e_{Z}).$
We use here that, with $X$ fixed, $\mathbb{E}_{W}[(W-e_{W}(X))^{2}]$ is the variance of a Bern$(e_{W}(X))$ random variable, so it equals $e_{W}(X)(1-e_{W}(X))$.
Finally, we make one more approximation to arrive at something computable by replacing the oracle functions with those estimated in the pilot study:
$V_{\mathrm{approx}}(e_{Z})$ $=n\bar{X}_{n}^{\top}\bigl{(}X^{\top}\mathrm{diag}(e_{W}(X)(1-e_{W}(X))X\bigr{)}^{-1}\bar{X}_{n}$
$\approx n\bar{X}_{n}^{\top}\bigl{(}X^{\top}\mathrm{diag}(\hat{e}_{W}(X)(1-\hat{e}_{W}(X))X\bigr{)}^{-1}\bar{X}_{n}$
$=\widehat{V}_{\mathrm{approx}}(e_{Z}).$
The function $\widehat{V}_{\mathrm{approx}}(e_{Z})$ is computable given just the new $X$ data and the functions estimated in the pilot study. Moreover, for a given sample, it is convex in the quantities $\{e_{Z}(X_{1}),\dots,e_{Z}(X_{n})\}$, so it can be efficiently minimized subject to any convex constraints. That is, with $\mathcal{C}_{n}\subseteq[0,1]^{n}$ being the feasible region for $e_{Z}$ imposed by our convex constraints, we can solve for
$\hat{e}_{Z}^{*}=\underset{e_{Z}\in\mathcal{C}_{n}}{\text{argmin }}\widehat{V}_{\mathrm{approx}}(e_{Z}).$ (11)
The two approximations that yield a tractable convex optimization problem are visualized in Figure 2, and we summarize our method in Algorithm 1.
Figure 2: Visualization of the approximations that move from $V$ to $\widehat{V}_{\mathrm{approx}}$. The left-right transition involves moving a matrix inverse out of an expectation to approximate the variance, and the top-bottom transition involves replacing oracle parameters with their estimates from the pilot study.
We could analogously define $e_{Z}^{*}$ as the design that minimizes the true variance $V(e_{Z})$ over $\mathcal{C}_{n}$, though there is no way to solve for this. In light of our approximations, a desirable property of our method is that, asymptotically, using $\hat{e}_{Z}^{*}$ recovers the optimal true variance $V(e_{Z}^{*})$ and the oracle estimate $\hat{\gamma}_{\mathrm{oracle}}$. The following theorem confirms that this is the case under certain assumptions.
Theorem 1.
Suppose that the following conditions all hold:
  1. 1.
    The pilot study has size $n_{0}=\Omega(n)$ that is not slower than linear in $n$.
  2. 2.
    The propensity scores are estimated sufficiently well:
    $\|\hat{e}_{W}(X)-e_{W}(X)\|_{\infty}=\underset{1\leq i\leq n}{\sup}\text{ }|\hat{e}_{W}(X_{i})-e_{W}(X_{i})|=o_{p}(1)$
    as $n\to\infty$.
  3. 3.
    The $m^{*}(X,Z,W)$ functionals are estimated sufficiently well:
    $\frac{1}{n}\|\hat{m}^{*}-m^{*}\|^{2}_{2}=\frac{1}{n}\sum_{i=1}^{n}(\hat{m}^{*}(X_{i},Z_{i},W_{i})-m^{*}(X_{i},Z_{i},W_{i}))^{2}=o_{p}(1),$
    as $n\to\infty$.
  4. 4.
    There exist $c,C>0$ and $n_{*}\geq d$ with $\|X_{i}\|^{2}_{2}\leq C$ for all $1\leq i\leq n$ and $X^{\top}X/n\succeq cI_{d}$ for all $n\geq n_{*}$, where $\succeq$ refers to the standard partial ordering on positive semi-definite matrices.
  5. 5.
    There exists $\eta>0$ such that $e_{W}(X)=p_{\mathrm{AT}}(X)+p_{\mathrm{C}}(X)e_{Z}(X)\in[\eta,1-\eta]$ for all $X\in\mathcal{X}$ and $e_{Z}\in\mathcal{C}_{n}$.
Let $e_{Z}^{*}$ and $\hat{e}_{Z}^{*}$ minimize the true variance $V(e_{Z})$ and estimated variance $\widehat{V}_{\mathrm{approx}}(e_{Z})$ respectively over $\mathcal{C}_{n}$. For given $(X,Z,W,Y)$ data, let $\hat{\gamma}_{\mathrm{oracle}}$ be the OLS estimate of $\gamma$ fit using oracle parameters as in (8), and let $\hat{\gamma}_{\mathrm{approx}}$ be the OLS estimate fit using approximations as in (9). Then
[…]
as $n\to\infty$.
The proof is in Appendix A.2. Theorem 1 confirms that the approximations made to arrive at a solvable optimization problem do no harm asymptotically. Assumption (1) guarantees that the size of the pilot study grows sufficiently with the overall sample size and can be accomplished simply by allotting $n_{0}=\alpha n$ for some fixed fraction $\alpha\in(0,1)$. Assumptions (2) and (3) give requirements about the quality of our oracle estimation, while Assumptions (4) and (5) ensure that certain quantities are bounded sufficiently away from zero. The first statement of Assumption (2) is implied if the analogous conditions hold for each of $\{\hat{p}_{\mathrm{C}}(X),\hat{p}_{\mathrm{AT}}(X),\hat{p}_{\mathrm{NT}}(X)\}$. Assumption (4) holds for free if one assumes that $p_{\mathrm{AT}}(X)$ and $p_{\mathrm{NT}}(X)$ are both uniformly bounded away from zero, or alternatively if one imposes an overlap constraint ensuring that every $e_{Z}\in\mathcal{C}_{n}$ is uniformly bounded away from zero and one.
Algorithm 1 Procedure to choose nudge probabilities
1: Input: pilot study $\mathcal{P}=\{X_{i},Z_{i},W_{i},Y_{i}\}_{1\leq i\leq n_{0}}$ obtained via $e_{Z}^{0}$, new data $X\in\mathbb{R}^{n\times d}$, feasible region $\mathcal{C}_{n}\subseteq[0,1]^{n}$ for $e_{Z}$ defined via convex constraints.
2:Using $\mathcal{P}$, construct estimates $\{\hat{p}_{\mathrm{C}}(X),\hat{p}_{\mathrm{AT}}(X),\hat{p}_{\mathrm{NT}}(X)\}$ of the compliance probabilities and $\hat{m}^{*}(X,Z,W)$ of the function $m^{*}(X,Z,W)$ defined in (6).
3:With $\hat{D}(e_{Z})=\hat{p}_{\mathrm{AT}}(X)+\hat{p}_{\mathrm{C}}(X)e_{Z}(X)$, define
4:  $\widehat{V}_{\mathrm{approx}}(e_{Z})=(X^{\top}\hat{D}(e_{Z})(1-\hat{D}(e_{Z}))X)^{-1}$
5:Solve the convex program (11) for $\hat{e}_{Z}^{*}$
6:Assign nudges with probabilities $e_{Z}^{*}$ to observe new data $\{Z_{i},W_{i},Y_{i}\}_{i\leq n}$.
7:With $\hat{D}(X,W)=\mathrm{diag}(W-(\hat{p}_{\mathrm{AT}}(X)+\hat{p}_{\mathrm{C}}(X)e_{Z}^{*}(X)))\in\mathbb{R}^{d\times d}$, compute
8:  $\hat{\gamma}_{\mathrm{approx}}=\bigl{(}X^{\top}\hat{D}(X,W)^{2}X\bigr{)}^{-1}X^{\top}\hat{D}(X,W)\bigl{(}Y-\hat{m}^{*}(X,Z,W)\bigr{)}\in\mathbb{R}^{d}$
9:
10:Output: $\hat{e}_{Z}^{*}$, $\hat{\gamma}_{\mathrm{approx}}$.
It is possible to generalize this procedure to heteroscedastic errors that depend on $X$ and $W$ by estimating the conditional variances $\sigma^{2}(X,W)=\text{Var}(\epsilon\text{ }|\text{ }X,W)$ in the pilot study. These can then be incorporated to change the second-stage regression to a weighted least squares procedure. For ease of exposition, we defer full details of this method to Appendix B. Of note, it yields a convex variance criterion, as in the homoscedastic case, provided that the variance ratio $\sigma^{2}(X,W=1)/\sigma^{2}(X,W=0)$ is always bounded in $[1/2,2]$ (see Theorem 3 in Appendix B).
3.4 Constraints on nudge assignment
A natural question to ask is what the optimal solution $\hat{e}_{Z}^{*}\in[0,1]^{n}$ looks like depending on the constraints imposed. This question has been considered for fully linear models without noncompliance by Li and Owen (2023a) and Morrison and Owen (2024). They prove that, when a monotonicity constraint is imposed that forces $e_{Z}(X_{1})\leq e_{Z}(X_{2})\leq\cdots\leq e_{Z}(X_{n})$, optimal designs consist of a few discrete levels of constant probability. Li and Owen (2023a) also find that optimal solutions have a few levels under only budget and gain constraints in a simple one-dimensional linear model.
As in those papers, we find that the RCT that has $e_{W}(X)=1/2$ everywhere is optimal, if attainable. Conceptually, this is because the $e_{W}(X)(1-e_{W}(X)$ entries in the diagonal matrix in $V_{\mathrm{approx}}(e_{Z})$ are all minimized at $e_{W}(X)=1/2$. Under noncompliance, this equates to choosing
$\hat{p}_{\mathrm{AT}}(X)+\hat{p}_{\mathrm{C}}(X)\hat{e}^{*}_{Z}(X)=\frac{1}{2}\implies\hat{e}_{Z}^{*}(X)=\frac{1-2\hat{p}_{\mathrm{AT}}(X)}{2\hat{p}_{\mathrm{C}}(X)}.$ (12)
In short, although we cannot choose $e_{W}(X)$ directly, to minimize variance we would choose $e_{Z}(X)$ so that we anticipate getting an RCT in the second stage based on our knowledge of the compliance probabilities.
The equation (12) implements this strategy by assigning lower nudge probability where the rate of being an always-taker is high and assigning higher nudge probability where the rate of being a never-taker is high. However, even in the unconstrained setting, this choice may be impossible if the (estimated) rate of noncompliance is too high, i.e., if
$\frac{1-2\hat{p}_{\mathrm{AT}}(X_{i})}{2\hat{p}_{\mathrm{C}}(X_{i})}\not\in[0,1]$
for some $X_{i}$. In this case, the solution is to threshold to whichever of 00 or $1$ is closer to this ratio. This choice only accounts for statistical considerations and may be infeasible on economic or ethical grounds.
Interestingly, a similar result holds under a budget constraint in the case where there is no noncompliance (or, more generally, the compliance probability $\hat{p}_{\mathrm{C}}(X_{i})$ is constant in $i$). That is, if we only impose that the average estimated treatment probability is some amount, i.e.,
$\bar{\hat{e}}_{W}=\frac{1}{n}\sum_{i=1}^{n}\hat{p}_{\mathrm{AT}}(X_{i})+\hat{p}_{\mathrm{C}}(X_{i})e_{Z}(X_{i})=\mu\in[0,1],$
so that a fixed fraction of subjects get the nudge in expectation, the optimal solution is the RCT with $\hat{e}_{W}(X)=\mu$. This result requires that $X$ have an intercept in its span. Interestingly, it also differs from the findings in Li and Owen (2023a) and Morrison and Owen (2024), whose objectives and models do not yield constant probabilities once a budget constraint has been added. This difference arises from special properties of the partially linear model and the particular objective function $\hat{V}_{\mathrm{approx}}(e_{Z})$ that it induces. We summarize all of this in the theorem below.
Theorem 2.
Consider the optimization problem
$\hat{e}_{Z}^{*}=\underset{e_{Z}\in\mathcal{C}_{n}}{\arg\min}\bar{X}_{n}^{\top}\widehat{V}_{\mathrm{approx}}(e_{Z})\bar{X}_{n},$
where $\mathcal{C}_{n}\subseteq[0,1]^{n}$ is a region defined by convex constraints.
(a) If $\mathcal{C}_{n}=[0,1]^{n}$ is unconstrained, then an optimal solution is to take
$\hat{e}_{Z}^{*}(X_{i})=\begin{cases}\frac{1-2\hat{p}_{\mathrm{AT}}(X_{i})}{2\hat{p}_{\mathrm{C}}(X_{i})},&\hat{p}_{\mathrm{AT}}(X_{i})\leq\frac{1}{2}\text{ and }\hat{p}_{\mathrm{AT}}(X_{i})+\hat{p}_{\mathrm{C}}(X_{i})\geq\frac{1}{2}\\ 0,&\hat{p}_{\mathrm{AT}}(X_{i})>\frac{1}{2}\\ 1,&\hat{p}_{\mathrm{AT}}(X_{i})+\hat{p}_{\mathrm{C}}(X_{i})<\frac{1}{2}.\end{cases}$
(b) If (i) $\mathcal{C}_{n}=\{e_{Z}\in[0,1]^{n}\text{ }|\text{ }\bar{e}_{W}=\mu\}$ for some $\mu\in[0,1]$ and (ii) the compliance probability $\hat{p}_{\mathrm{C}}(X)$ is constant, and (iii) $1_{n}\in\mathrm{span}(X)$, then an optimal solution is to take
$\hat{e}_{Z}^{*}(X_{i})=\frac{\mu-\bar{\hat{p}}_{\mathrm{AT}}}{\hat{p}_{\mathrm{C}}},$
where $\bar{\hat{p}}_{\mathrm{AT}}$ is the sample average of the always-taker probabilities, provided that this fraction is in $[0,1]$.
Theorem 2 provides intuition about the nature of optimal solutions under minimal constraints. In the unconstrained case, an increase in $\hat{p}_{\mathrm{AT}}(X_{i})$ coincides with a decrease in $\hat{e}_{Z}^{*}(X_{i})$ to try to balance the number of treated and control units, for instance.
4 Triage example
In this section, we outline an example application of this procedure that is motivated by the problem of patient triage in emergency departments (EDs). When a patient arrives in the ED, they are assigned a level of care based on their perceived need and bed availability, with the highest level of care being an intensive care unit (ICU). Because ICUs require greater resources and are constrained in their bed and nursing availability, a natural question of interest is the causal effect of being sent to the ICU on a patient’s health outcomes. This question could impact decisions about both the size of the ICU (Phua et al., 2020) and future triaging, which may be able to exploit the heterogeneous effects of treatment (Chang et al., 2017).
Because full randomization would be highly unethical, we consider a semi-synthetic example based on a study design in which the experimenter assigns a given patient to either the ICU ($Z=1$) or less intensive care ($Z=0$) and their doctor can override if deemed necessary. This example thus falls naturally into our framework because the experimenter’s proposed triage is a nudge $Z$ that can depend on the patient’s observed covariates $X$ and that need not agree with the final triage assignment $W$. Moreover, it is natural to expect that patients whose encouragement is overridden by the physician likely differ in some unmeasured way, introducing confoundedness.
We use a de-identifed electronic health records dataset that consists of $7281$ adult inpatient admissions at an academic hospital and a trauma center between 2015 and 2019. This dataset includes each patient’s length of stay in the ED. We take this to be our response variable of interest, with a lower length of stay indicating a preferred outcome. It also includes several covariates about the patient, including their demographic information and various vital signs.
To these variables, we add the risk score developed by Nguyen et al. (2021). This is a machine learning-based prediction of the patient’s probability of going to the ICU based on their covariates. In our motivating setting, this would be a natural variable to use to constrain the experimental design, with high-risk patients having a higher probability of encouragement than low-risk patients. For this reason, we return to it when imposing constraints later in our analysis.
Because our dataset is observational and does not come from the proposed study design, we do not have access to genuine noncompliance data, and our example is necessarily semi-synthetic. Nonetheless, we use the observed data to fit sensible models for $Y$ given $X$ and $W$, which we take as a starting point for our analysis. Specifically, we proceed as follows:
  1. 1.
    Use the non-ICU data ($W=0$) to fit a flexible baseline model $Y\sim f(X)$ via a random forest. Here, $Y\in\mathbb{R}^{n}$ is the length of stay (in hours) for $n$ patients and $X\in\mathbb{R}^{n\times 7}$ is a dataset containing a subset of covariates deemed particularly important. These are: hours in the ED prior to triage, weight, age, systolic blood pressure (SBP), pulse, risk score, and an intercept. One could also use a much richer class of covariates in this step and restrict to fewer for the linear model to follow, but for simplicity we opted to keep the dimensions the same throughout.
  2. 2.
    Use the ICU data ($W=1$) to fit a regression model $Y-f(X)\sim WX\gamma$, where $X\in\mathbb{R}^{n\times 7}$ is as above.
  3. 3.
    Choose functions $\{p_{\mathrm{C}}(X),p_{\mathrm{AT}}(X),p_{\mathrm{NT}}(X)\}$ for the noncompliance probabilities and assign compliance classes to patients via these probabilities. These are entirely synthetic, since there were no nudges and hence no notion of noncompliance in the data. For simplicity and interpretability, we choose these to depend only on the risk score of Nguyen et al. (2021). We base our choice on the heuristic that it is natural to expect higher rates of noncompliance at extremes of this risk score. Full formulas are available in Appendix C.
  4. 4.
    Choose a distribution for the error term $\varepsilon$ that depends on compliance class. We take $\varepsilon$ to include additive terms for each noncompliance class, i.e.,
    $\varepsilon_{i}\overset{\mathrm{ind}}{\sim}\mathds{1}\{\text{AT}_{i}\}\gamma_{\mathrm{AT}}+\mathds{1}\{\text{NT}_{i}\}\gamma_{\mathrm{NT}}+N(0,\sigma^{2}).$
    The particular coefficients $\gamma_{\mathrm{AT}}$ and $\gamma_{\mathrm{NT}}$ are free parameters that tune how much the noncompliance affects outcomes. In addition, $\sigma^{2}$ is a noise parameter that can be tuned.
Further details about the simulation procedure are available in Appendix C.
Figure 3: Results of our convex optimization procedure under an array of constraint settings: (a) unconstrained, (b) monotonicity constraint, (c) budget constraint, (d) budget and monotonicity constraints (e) budget, monotonicity, and low gain constraint, and (f) budget, monotonicity, and high gain constraint. All solutions besides (a) are obtained via the CVXR package (Fu et al., 2020).
In Figure 3, we present the optimal solution $\hat{e}_{Z}^{*}$ under an array of constraint settings interpreted below. We used $n=6000$. The risk scores in that figure have been rank transformed to have a uniform distribution. The constraints imposed, in various combinations, are:
  1. 1.
    No constraints, i.e., $\mathcal{C}_{n}=[0,1]^{n}$. Here, the solution is known perfectly via Theorem 2 and convex optimization is not needed.
  2. 2.
    Monotonicity constraint that $\hat{e}_{Z}(X)$ is nondecreasing in the risk score $r(X)$. This serves as a fairness constraint to force the nudge propensity to be no lower for a higher risk patient than for a lower risk one.
  3. 3.
    Budget constraint with $\mathcal{C}_{n}=\{e_{Z}\in[0,1]^{n}\text{ }|\text{ }\overline{p_{\mathrm{AT}}}+\overline{p_{C}e_{Z}}=0.3\}$. Under this constraint the expected proportion of subjects treated is $30\%$.
  4. 4.
    Gain constraint that
    $\sum_{i=1}^{n}e_{Z}(X_{i})r(X_{i})\geq\rho\sum_{i=k}^{n}r(X_{(i)}),$
    where $\rho$ is a free parameter and the parenthetical subscript on the right-hand side indicates that we sort risk scores from smallest to largest. In words, the expected total risk among those encouraged is at least some fraction $\rho$ times what it would be under a regression discontinuity design (RDD) in which the encouragement is given deterministically to the patients with the highest risk score. The lower summand $k$ on the right-hand side is dictated by the other constraints of the problem; for example, under a budget constraint, $k$ would be such that the average encouragement probability over the whole patient population is as stipulated.
    This constraint simulates one that might arise in a real setting, in which there is some expected benefit to be derived by giving the encouragement to particular units over others. Here, the risk score is a proxy for such benefit. If this $\rho$ parameter were brought closer to $1$, the result would tend to a regression discontinuity design in which precisely the top patients by risk score were given the encouragement.
Figure 3 shows the graphs for optimal solutions in these settings. The unconstrained solution is clearly inappropriate in a real ICU, since it assigns the highest nudge probabilities to cases with the lowest risk scores and vice versa. This optimizes statistical efficiency by assigning more encouragements to patients more likely to be never-takers and fewer to patients more likely to be always-takers in order to arrive at something closer to an RCT. However it is completely inappropriate, recommending the ICU for those who need it least or not at all, while withholding that recommendation for those with the greatest need. It also does not take account of the limited capacity of the ICU. Adding a monotonicity constraint flattens and reverses the curve. The large change in $e_{Z}$ at the extreme ends of the risk score in this solution appears to be an artifact of CVX; we observed that a flat solution at $e_{Z}=0.49$ yields a slightly smaller objective value. The budget constraint, which allows only $30\%$ expected treatment, depresses encouragement probabilities, though the benefit remains to encourage lower risk patients more often to combat the fact that there are more never-takers in that regime. Finally, the bottom two panels add, respectively, a “low” gain constraint with $\rho=0.7$ and a “high” gain constraint with $\rho=0.95$. As anticipated, the solution moves closer to an RDD as $\rho$ tends to one. We also compare the values of the objective for each solution in Appendix C, finding that the constrained problems have objective values between $7\%$ and $83\%$ larger than the unconstrained one.
Figure 4: Variance ratio (left) and MSE ratio (right) of our proposed method to the baseline RDD. The dashed red line at $1$ corresponds to performance equal to the baseline.
Separately, we assess the variance and mean-squared error (MSE) of our procedure to estimate $\tau_{\mathrm{LATE}}$. We do so under both a budget constraint that $\overline{\hat{e}}_{W}=0.4$ and a monotonicity constraint to capture bed availability and fairness constraints that might naturally arise. As a baseline, we compare our optimal design $\hat{e}_{Z}^{*}$ to the RDD that gives encouragement deterministically to the highest risk patients, with the cutoff determined by satisfying the same $40\%$ budget constraint. This RDD is a natural comparison point that one might default to in a hospital if not using our method. At each of $n\in\{250,500,1000,2000,3000,4000,5000,6000\}$, we compute $\hat{\tau}_{\mathrm{LATE}}$ under both our design and the baseline design. We repeat this over $100$ simulations. Finally, we compute the sample variance and MSE of each procedure and compare.
These results are presented in Figure 4, which plots the variance ratios and MSE ratios of our method to the baseline RDD. The latter is in the denominator of each ratio so that values below one indicate improvements. We see that our method consistently outperforms the baseline RDD in both variance and MSE across all values of $n$ tested.
The results in Figure 4 suggest that our method is capable of producing meaningful gains over less thoughtful designs. Moreover, when more constraints are added, it becomes less clear what a baseline design choice should even be, since identifying a feasible design becomes non-obvious.
Acknowledgments
The authors thank John Cherian, Isaac Gibbs, Samir Khan, Harrison Li, and Anav Sood for helpful discussions and comments and thank Jonathan H. Chen for providing data resources that greatly enriched this project. T. M. was partly supported by a B. C. and E. J. Eaves Stanford Graduate Fellowship. M. N. was supported by the Stanford DARE fellowship. A. O. was supported by grant DMS-2152780 from the U.S. National Science Foundation.
Appendix A Proofs
A.1 Proof of LATE identifiability
Here, we reprove Equation (2) for completeness. We have
$\tau_{\mathrm{LATE}}$ $=\mathbb{E}_{X}\left[\frac{\mathbb{E}[Y\text{ }|\text{ }X,Z=1]-\mathbb{E}[Y\text{ }|\text{ }X,Z=0]}{\mathbb{E}[W\text{ }|\text{ }X,Z=1]-\mathbb{E}[W\text{ }|\text{ }X,Z=0]}\right]$
[…]
$=\mathbb{E}[Y(1)-Y(0)\text{ }|\text{ }\text{complier}].$
The penultimate line uses Assumption 1 that there are no defiers, so the only units for which $Y(W(1))\neq Y(W(0))$ are compliers. Exclusion and conditional identifiability come from Definition 1 for our instrumental variable. We did not need to use the relevance condition. The final line is simply Bayes’ theorem. We used Assumption 2 throughout to keep $p_{\mathrm{C}}(X)>0$.
A.2 Proof of Theorem 1
In light of the approximations in Figure 2, we begin with two lemmas showing that
(1) $\underset{e_{Z}\in\mathcal{C}_{n}}{\sup}\text{ }|V(e_{Z})-V_{\mathrm{approx}}(e_{Z})|\to 0$,  and
[…]
Lemma 1.
Under the conditions of Theorem 1,
$\underset{e_{Z}\in\mathcal{C}_{n}}{\sup}\text{ }|V(e_{Z})-V_{\mathrm{approx}}(e_{Z})|=O(n^{-1/2})$
as $n\to\infty$, where the implied constant depends only on $\eta$, $c$ and $C$.
Proof.
Recall that $V(e_{Z})$ is the conditional variance of $\sqrt{n}\bar{X}_{n}^{\top}\hat{\gamma}_{*}$ given $X_{1},\dots,X_{n}$. Fix $e_{Z}\in\mathcal{C}_{n}$, and define $e_{W}(X)=p_{\mathrm{AT}}(X)+p_{\mathrm{C}}(X)e_{Z}(X)$. Let $D=\mathrm{diag}(W_{i}-e_{W}(X_{i}))$. Then $W_{i}\stackrel{{\scriptstyle\mathrm{ind}}}{{\sim}}\mathrm{Bern}(e_{W}(X_{i}))$ and we will show that
$n\bigl{|}\bar{X}_{n}^{\top}\mathbb{E}[(X^{\top}D^{2}X)^{-1}]\bar{X}_{n}-\bar{X}_{n}^{\top}\mathbb{E}[X^{\top}D^{2}X]^{-1}\bar{X}_{n}\bigr{|}=O(n^{-1/2})$
for any vector of nudge propensities $e_{Z}\in\mathcal{C}_{n}$.
By the Cauchy-Schwarz inequality and the assumption that $\|X_{i}\|_{2}^{2}\leq C$,
$n\bigl{|}\,\bar{X}_{n}^{\top}\mathbb{E}[(X^{\top}D^{2}X)^{-1}]\bar{X}_{n}-\bar{X}_{n}^{\top}\mathbb{E}[X^{\top}D^{2}X]^{-1}\bar{X}_{n}\bigr{|}\\ \leq Cn\big{\|}\mathbb{E}[(X^{\top}D^{2}X)^{-1}]-\mathbb{E}[X^{\top}D^{2}X]^{-1}\big{\|}_{\mathrm{op}},$ (13)
where $\|\cdot\|_{\mathrm{op}}$ denotes the operator norm of a matrix. We will show that the right-hand side of (13) converges to zero. To prove this, we proceed in steps. First, write
$X^{\top}D^{2}X=\sum_{i=1}^{n}(W_{i}-\hat{e}_{W}(X_{i}))^{2}X_{i}X_{i}^{\top}.$
Define the random matrices
$M_{i}=(W_{i}-e_{W}(X_{i}))^{2}X_{i}X_{i}^{\top}-e_{W}(X_{i})(1-e_{W}(X_{i}))X_{i}X_{i}^{\top},$
which each have mean zero and are independent. Moreover,
$\|M_{i}^{2}\|_{\mathrm{op}}\leq\|(X_{i}X_{i}^{\top})^{2}\|_{\mathrm{op}}\leq C^{2}$
by assumption that $\|X_{i}\|_{2}^{2}\leq C$. By the matrix Hoeffding inequality (see Tropp (2011)), for any $t>0$:
$\mathbb{P}\left(\|X^{\top}D^{2}X-\mathbb{E}[X^{\top}D^{2}X]\|_{\mathrm{op}}\geq t\right)$ ($\leq\mathbb{P}\left(\left\{\lambda_{\max}\left(\sum_{i=1}^{n}M_{i}\right)\geq t\right\}\right)$
+($\quad+\mathbb{P}\left(\left\{\lambda_{\min}\left(\sum_{i=1}^{n}M_{i}\right)\leq-t\right\}\right)$
$\leq 2de^{-t^{2}/8nC^{2}}.$ (14)
Now let $M=X^{\top}D^{2}X$. Then
$\|n\mathbb{E}[M^{-1}]-n\mathbb{E}[M]^{-1}\|_{\mathrm{op}}$ $=n\bigl{\|}\mathbb{E}[M^{-1}-\mathbb{E}[M]^{-1}]\bigr{\|}_{\mathrm{op}}$
$\leq n\mathbb{E}\bigl{[}\|M^{-1}-\mathbb{E}[M]^{-1}\|_{\mathrm{op}}\bigr{]}$
$=n\mathbb{E}\bigl{[}\bigl{\|}M^{-1}\bigl{(}\mathbb{E}[M]-M\bigr{)}\mathbb{E}[M]^{-1}\bigr{\|}_{\mathrm{op}}\bigr{]}$
$\leq n\mathbb{E}\bigl{[}\|M^{-1}\|_{\mathrm{op}}\times\|M-\mathbb{E}[M]\|_{\mathrm{op}}\bigr{]}\times\bigl{\|}\mathbb{E}[M]^{-1}\bigr{\|}_{\mathrm{op}}.$
Now, we can use our assumptions (3) and (4) to see that, for $n$ sufficiently large,
$\mathbb{E}[M]$ $=\sum_{i=1}^{n}e_{W}(X_{i})(1-e_{W}(X_{i}))X_{i}X_{i}^{\top}$
$\succeq\sum_{i=1}^{n}\eta(1-\eta)X_{i}X_{i}^{\top}$
$\succeq n\eta(1-\eta)cI_{d}.$
Hence, $\|\mathbb{E}[M]^{-1}\|_{\mathrm{op}}=O(n^{-1})$. Similarly, for $n$ sufficiently large,
$M$ $=\sum_{i=1}^{n}\bigl{(}W-e_{W}(X_{i})\bigr{)}^{2}X_{i}X_{i}^{\top}\succeq\sum_{i=1}^{n}\eta^{2}X_{i}X_{i}^{\top}\succeq n\eta^{2}cI_{d}$
and so $\|M^{-1}\|_{\mathrm{op}}=O(n^{-1})$. As for the last term, we can use the matrix Hoeffding result (A.2) to get
$\mathbb{E}\bigl{[}\|M-\mathbb{E}[M]\|_{\mathrm{op}}\bigr{]}$ $=\int_{0}^{\infty}\mathbb{P}\bigl{(}\|M-\mathbb{E}[M]\|_{\mathrm{op}}\geq t\bigr{)}\,\mathrm{d}t$
$\leq 2d\int_{0}^{\infty}e^{-t^{2}/8nC^{2}}\,\mathrm{d}t$
$=d\sqrt{8\pi nC^{2}}.$
In summary,
$|V(e_{Z})-V_{\mathrm{approx}}(e_{Z})|$ $\leq nC\underbrace{\bigl{\|}\mathbb{E}[M]^{-1}\bigr{\|}_{\mathrm{op}}}_{O(n^{-1})}\mathbb{E}\bigl{[}\underbrace{\|M^{-1}\|_{\mathrm{op}}}_{O(n^{-1})}\|M-\mathbb{E}[M]\|_{\mathrm{op}}\bigr{]}$
$=O(n^{-1})\underbrace{\mathbb{E}\bigl{[}\|M-\mathbb{E}[M]\|_{\mathrm{op}}\bigr{]}}_{O(n^{1/2})}$
$=O(n^{-1/2}).$
Thus, we have an upper bound of order $O(n^{-1/2})$ that does not depend on our choice of $e_{Z}\in\mathcal{C}_{n}$, so the limit is indeed zero uniformly. ∎
Lemma 2.
Under the conditions of Theorem 1,
[…]
as $n\to\infty$.
Proof.
Let $E=\mathrm{diag}(e_{W}(X)(1-e_{W}(X))$ and $\hat{E}=\mathrm{diag}(\hat{e}_{W}(X)(1-\hat{e}_{W}(X))$. As in Lemma 1, it suffices to show that
[…]
Write $Q=X^{\top}EX$ and $\hat{Q}=X^{\top}\hat{E}X$, in which case
$n\|Q^{-1}-\hat{Q}^{-1}\|_{\mathrm{op}}$ $=n\|Q^{-1}(\hat{Q}-Q)\hat{Q}^{-1}\|_{\mathrm{op}}$
$\leq n\|Q^{-1}\|_{\mathrm{op}}\|\hat{Q}^{-1}\|_{\mathrm{op}}\|Q-\hat{Q}\|_{\mathrm{op}}.$
The first operator norm term is $O(n^{-1})$ as shown in Lemma 1. For the second, recall the assumption that $\|\hat{e}_{W}(X)-e_{W}(X)\|_{\infty}$ is $o_{p}(1)$ as $n\to\infty$. With probability tending to one, therefore, $\hat{e}_{W}(X)(1-\hat{e}_{W}(X))$ is within $\eta(1-\eta)/2$ of the true $e_{W}(X)(1-e_{W}(X))$ for all $X$. Hence, with probability tending to one,
$\hat{Q}$ $=\sum_{i=1}^{n}\hat{e}_{W}(X_{i})(1-\hat{e}_{W}(X_{i}))X_{i}X_{i}^{\top}$
$\succeq\sum_{i=1}^{n}\frac{1}{2}\eta(1-\eta)X_{i}X_{i}^{\top}$
$\succeq\frac{n}{2}\eta(1-\eta)cI_{d},$
meaning that $\|\hat{Q}^{-1}\|_{\mathrm{op}}=O_{p}(n^{-1})$.
Finally, to address the last term, we have
$\|Q-\hat{Q}\|_{\mathrm{op}}$ $=\|X^{\top}(E-\hat{E})X\|_{\mathrm{op}}$
$=\underset{\|v\|_{2}=1}{\sup}\sum_{i=1}^{n}(X_{i}^{\top}v)^{2}(E_{ii}-\hat{E}_{ii})$
$\leq\underset{\|v\|_{2}=1}{\sup}\sqrt{\sum_{i=1}^{n}(X_{i}^{\top}v)^{4}}\sqrt{\sum_{i=1}^{n}(E_{ii}-\hat{E}_{ii})^{2}}.$
Now $\sqrt{\sum_{i=1}^{n}(X_{i}^{\top}v)^{4}}=O(n^{1/2})$ because $\|X_{i}\|_{2}\leq C$, and
$\sum_{i=1}^{n}(E_{ii}-\hat{E}_{ii})^{2}$ $=\sum_{i=1}^{n}\bigl{(}e_{W}(X_{i})-\hat{e}_{W}(X_{i})\bigr{)}^{2}\bigl{(}1+e_{W}(X_{i})+\hat{e}_{W}(X_{i})\bigr{)}^{2}$
$\leq 9\sum_{i=1}^{n}(e_{W}(X_{i})-\hat{e}_{W}(X_{i}))^{2}$
$=o_{p}(n)$
using $1+e_{W}(X_{i})+\hat{e}_{W}(X_{i})\leq 3$ and $\|\hat{e}_{W}-e_{W}\|_{\ell^{\infty}}=o_{p}(1)$. As a result, $\|Q-\hat{Q}\|_{\mathrm{op}}=o_{p}(n)$ and then
[…]
$V(\hat{e}_{Z}^{*})-V(e_{Z}^{*})$ $=V(\hat{e}_{Z}^{*})-V_{\mathrm{approx}}(\hat{e}_{Z}^{*})$
$\quad+V_{\mathrm{approx}}(\hat{e}_{Z}^{*})-V_{\mathrm{approx}}(e_{Z}^{*})$
$\quad+V_{\mathrm{approx}}(e_{Z}^{*})-V(e_{Z}^{*}).$
[…]
$\hat{\gamma}_{\mathrm{oracle}}$ $=(X^{\top}D^{2}X)^{-1}X^{\top}D(Y-m^{*}),\quad\text{and}$
$\quad\hat{\gamma}_{\mathrm{approx}}$ $=(X^{\top}\hat{D}^{2}X)^{-1}X^{\top}\hat{D}(Y-\hat{m}^{*}).$
As in Lemma 1, $D=\mathrm{diag}(W_{i}-e_{W}(X_{i}))$, and $\hat{D}=\mathrm{diag}(W_{i}-\hat{e}_{W}(X_{i}))$, while $m^{*}=m^{*}(X,Z,W)$ and $\hat{m}^{*}=\hat{m}^{*}(X,Z,W)$ are $n$-dimensional vectors with entries as defined in (6) and (10), respectively. Define $M=X^{\top}D^{2}X$ and $\hat{M}=X^{\top}\hat{D}^{2}X$. Then
$\|\hat{\gamma}_{\mathrm{oracle}}-\hat{\gamma}_{\mathrm{approx}}\|_{2}$ $=\|M^{-1}X^{\top}D(Y-m^{*})-\hat{M}^{-1}X^{\top}\hat{D}(Y-\hat{m}^{*})\|_{2}$
$\leq\|M^{-1}X^{\top}D(Y-m^{*})-M^{-1}X^{\top}D(Y-\hat{m}^{*})\|_{2}$
$\phantom{=}+\|M^{-1}X^{\top}D(Y-\hat{m}^{*})-M^{-1}X^{\top}\hat{D}(Y-\hat{m}^{*})\|_{2}$
$\phantom{=}+\|M^{-1}X^{\top}\hat{D}(Y-\hat{m}^{*})-\hat{M}^{-1}X^{\top}\hat{D}(Y-\hat{m}^{*})\|_{2}$
$:=T_{1}+T_{2}+T_{3}.$
It suffices to show that each of these three terms converges in probability to zero.
For term $T_{1}$, we have
$T_{1}$ $=\|M^{-1}X^{\top}D(\hat{m}^{*}-m^{*})\|_{2}\leq\underbrace{\|M^{-1}\|_{\mathrm{op}}}_{O(n^{-1})}\underbrace{\|X^{\top}D\|_{\mathrm{op}}}_{O(n^{1/2})}\underbrace{\|\hat{m}^{*}-m^{*}\|_{2}}_{o_{p}(n^{1/2})}$
$=o_{p}(1).$
Here, we use the assumption in the theorem statement that $\|\hat{m}^{*}-m^{*}\|_{2}=o_{p}(n^{1/2})$.
For term $T_{2}$, we have
$T_{2}$ $=\|M^{-1}X^{\top}(D-\hat{D})(Y-\hat{m}^{*})\|_{2}$
$\leq\|M^{-1}\|_{\mathrm{op}}\|X^{\top}\|_{\mathrm{op}}\|D-\hat{D}\|_{\mathrm{op}}\|Y-\hat{m}^{*}\|_{2}.$
Define $\tilde{\varepsilon}(X,W)=\varepsilon-\mathbb{E}[\varepsilon\text{ }|\text{ }X,W]$. We have
$\|Y-\hat{m}^{*}\|_{2}$ $\leq\|Y-m^{*}\|_{2}+\|m^{*}-\hat{m}^{*}\|_{2}$
$=\|(W-e_{W}(X))X\gamma+\tilde{\varepsilon}\|_{2}+\|M^{*}-\hat{m}^{*}\|_{2}$
$\leq\underbrace{\|X\gamma\|_{2}}_{O(n^{1/2})}+\underbrace{\|\tilde{\varepsilon}\|_{2}}_{O_{p}(n^{1/2})}+\underbrace{\|M^{*}-\hat{m}^{*}\|_{2}}_{o_{p}(n^{1/2})}.$
These bounds follow respectively from the assumption that $\|X_{i}\|_{2}^{2}\leq C$, the assumption that the $\tilde{\varepsilon}$ terms are independent with uniformly bounded variance, and the assumption that $\|M^{*}-\hat{m}^{*}\|_{2}=o_{p}(n^{1/2})$. This sum is therefore $O_{p}(n^{1/2})$. Then
$T_{2}\leq\underbrace{\|M^{-1}\|_{\mathrm{op}}}_{O(n^{-1})}\underbrace{\|X^{\top}\|_{\mathrm{op}}}_{O(n^{1/2})}\underbrace{\|D-\hat{D}\|_{\mathrm{op}}}_{o_{p}(1)}\underbrace{\|Y-\hat{m}^{*}\|_{2}}_{O_{p}(n^{1/2})}=o_{p}(1).$
Finally,
$T_{3}$ $=\|(M^{-1}-\hat{M}^{-1})X^{\top}\hat{D}(Y-\hat{m}^{*})\|$
$\leq\|M^{-1}-\hat{M}^{-1}\|_{\mathrm{op}}\|X^{\top}\hat{D}\|_{\mathrm{op}}\|Y-\hat{m}^{*}\|_{2}.$
We have shown that the latter two terms are $O(n^{1/2})$ and $O_{p}(n^{1/2})$ already, so it suffices to show that the first term is $o_{p}(n^{-1})$. Indeed,
$\|M^{-1}-\hat{M}^{-1}\|_{\mathrm{op}}$ $\leq\|M^{-1}\|_{\mathrm{op}}\|\hat{M}^{-1}\|_{\mathrm{op}}\|M-\hat{M}^{-1}\|_{\mathrm{op}}$
$=\|M^{-1}\|_{\mathrm{op}}\|\hat{M}^{-1}\|_{\mathrm{op}}\|X^{\top}(D-\hat{D})X\|_{\mathrm{op}}$
$\leq\underbrace{\|M^{-1}\|_{\mathrm{op}}}_{O(n^{-1})}\underbrace{\|\hat{M}^{-1}\|_{\mathrm{op}}}_{O_{p}(n^{-1})}\underbrace{\|X\|_{\mathrm{op}}^{2}}_{O(n)}\underbrace{\|D-\hat{D}\|_{\mathrm{op}}}_{o_{p}(1)}.$
This product is thus $o_{p}(n^{-1})$, which completes the proof.
A.3 Proof of Theorem 2
For part (a) of the theorem, let $f(e_{Z})$ be the objective function, i.e.,
$f(e_{Z})$ $=\bar{X}_{n}^{\top}\widehat{V}_{\mathrm{approx}}(e_{Z})\bar{X}_{n}$
$=\bar{X}_{n}^{\top}\bigl{(}X^{\top}\mathrm{diag}(\hat{e}_{W}(X)(1-\hat{e}_{W}(X)))X\bigr{)}^{-1}\bar{X}_{n},$
where $\hat{e}_{W}(X)=\hat{p}_{\mathrm{AT}}(X)+\hat{p}_{\mathrm{C}}(X)e_{Z}(X)$. Clearly, if for candidate probability vectors $e_{Z,1}$ and $e_{Z,2}$, $\widehat{V}_{\mathrm{approx}}(e_{Z,1})\preceq\widehat{V}_{\mathrm{approx}}(e_{Z,2})$, then $f(e_{Z,1})\leq f(e_{Z,2})$. Hence,
$\hat{e}_{W,1}(1-\hat{e}_{W,1})\succeq\hat{e}_{W,2}(1-\hat{e}_{W,2})\implies f(e_{Z,1})\leq f(e_{Z,2}).$
Since $x(1-x)$ is maximized at $x=1/2$, the optimal solution if possible is to take
$\hat{e}_{W}(X_{i})=\frac{1}{2}\iff e_{Z}(X_{i})=\frac{1-2\hat{p}_{\mathrm{AT}}(X_{i})}{2\hat{p}_{\mathrm{C}}(X_{i})}.$ (15)
If this is not in $[0,1]$, then the optimal solution is to take either $e_{Z}(X_{i})=0$ or $e_{Z}(X_{i})=1$, depending on which yields $\hat{e}_{W}(X_{i})$ closer to $1/2$. From (15), we see that this will be 00 if $\hat{p}_{\mathrm{AT}}>1/2$ and $1$ if $\hat{p}_{\mathrm{AT}}+\hat{p}_{\mathrm{C}}<1/2$. In summary,
$\hat{e}_{Z}^{*}(X_{i})=\begin{cases}\frac{1-2\hat{p}_{\mathrm{AT}}(X_{i})}{2\hat{p}_{\mathrm{C}}(X_{i})},&p_{\mathrm{AT}}\leq\frac{1}{2}\text{ and }p_{\mathrm{AT}}+p_{\mathrm{C}}\geq\frac{1}{2}\\ 0,&p_{\mathrm{AT}}>\frac{1}{2}\\ 1,&p_{\mathrm{AT}}+p_{\mathrm{C}}<\frac{1}{2}.\end{cases}$
For part (b), consider the case where we have a budget constraint, i.e., $\mathcal{C}_{n}=\{e_{Z}\in[0,1]^{n}\text{ }|\text{ }\bar{\hat{e}}_{W}=\mu\}$. Taking $e=e_{Z}$, the Lagrangian is
$L(e;\lambda,\tilde{\lambda},\nu)$ $=\bar{X}_{n}^{\top}\bigl{(}X^{\top}\mathrm{diag}(\hat{e}_{W}(X)(1-\hat{e}_{W}(X)))X\bigr{)}^{-1}\bar{X}_{n}$
$+\sum_{i=1}^{n}\lambda_{i}(e_{i}-1)-\tilde{\lambda}_{i}e_{i}+\nu\bigl{(}(\hat{p}_{\mathrm{AT}}(X_{i})+\hat{p}_{\mathrm{C}}(X_{i})e_{i})^{\top}1-\mu\bigr{)},$
where $\lambda_{i}$ and $\tilde{\lambda}_{i}$ are Lagrange multipliers for the constraints $e_{i}\leq 1$ and $e_{i}\geq 0$, while $\nu$ is the Lagrange multiplier for the budget constraint. Let $\hat{D}(e)=\mathrm{diag}(\hat{e}_{W}(X)(1-\hat{e}_{W}(X))$ and $\hat{M}(e)=X^{\top}\hat{D}(e)X$, in which case
$\frac{\partial}{\partial e_{i}}\bar{X}_{n}^{\top}\hat{M}(e)^{-1}\bar{X}_{n}$ $=\bar{X}_{n}^{\top}\left(\frac{\partial}{\partial e_{i}}\hat{M}(e)^{-1}\right)\bar{X}_{n}$
$=-\bar{X}_{n}^{\top}\hat{M}(e)^{-1}\left(\frac{\partial}{\partial e_{i}}\hat{M}(e)\right)\hat{M}(e)^{-1}\bar{X}_{n}$
$=-\bar{X}_{n}^{\top}\hat{M}(e)^{-1}X^{\top}\hat{D}_{i}(e)X\hat{M}(e)^{-1}\bar{X}_{n},$
where
$\hat{D}_{i}(e)=\frac{\partial}{\partial e_{i}}\hat{D}(e)=\begin{bmatrix}0\\ &\ddots\\ &&\hat{p}_{\mathrm{C}}(X_{i})(1-2\hat{e}_{W}(X_{i}))\\ &&&\ddots\\ &&&&0\end{bmatrix}.$
The stationarity condition for the Lagrangian is then
$\frac{\partial L}{\partial e_{i}}=-\bar{X}_{n}^{\top}\hat{M}(e)^{-1}X^{\top}\hat{D}_{i}(e)X\hat{M}(e)^{-1}\bar{X}_{n}+\lambda_{i}-\tilde{\lambda}_{i}-\nu=0\text{ for all }1\leq i\leq n.$
To prove that $\hat{e}_{Z}^{*}$ should be constant, first note that this would imply that $\lambda_{i}=\tilde{\lambda}_{i}=0$ for all $i$ by complementary slackness. This would then reduce to
$\bar{X}_{n}^{\top}\hat{M}(e)^{-1}X^{\top}\hat{D}_{i}(e)X\hat{M}(e)^{-1}\bar{X}_{n}=\nu\text{ for all }1\leq i\leq n.$
Thus, it suffices to show that, for $\hat{e}_{Z}$ chosen such that $\hat{e}_{W}=\mu 1_{n}$ is a constant vector, the gradient of the objective is the same in each entry. At this choice, $D(\hat{e}_{Z})=\mu(1-\mu)I_{n}$ is a multiple of the identity. Moreover, note that $\bar{X}_{n}=X^{\top}1_{n}$. Then
$\bar{X}_{n}^{\top}\hat{M}(e)^{-1}X^{\top}\hat{D}_{i}(e)X\hat{M}(e)^{-1}\bar{X}_{n}$ $\propto 1_{n}^{\top}X(X^{\top}X)^{-1}X^{\top}\hat{D}_{i}(e)X(X^{\top}X)^{-1}X^{\top}1_{n}$
$=1_{n}^{\top}H\hat{D}_{i}(e)H1_{n}$
$=1_{n}^{\top}\hat{D}_{i}(e)1_{n}$
$=\hat{p}_{\mathrm{C}}(X_{i})(1-2\mu).$
Here, $H=X(X^{\top}X)^{-1}X^{\top}$ is the standard hat matrix, and we use that $H1_{n}=1_{n}$ since $1_{n}\in\mathrm{span}(X)$ by assumption. Moreover, since the compliance probability $\hat{p}_{\mathrm{C}}$ is constant by assumption, this quantity is constant in $i$, and so the KKT conditions are satisfied. Consequently, $\hat{e}_{Z}^{*}=(\mu-\bar{\hat{p}}_{\mathrm{AT}})/{\hat{p}_{\mathrm{C}}}$ is a solution to the constrained optimization.
Appendix B Generalization to weighted least squares
Suppose first that we allow $\text{Var}(\varepsilon\text{ }|\text{ }X,Z,W)=\sigma^{2}(X)$ to vary with $X$ (we explain subsequently the case of dependence on both $X$ and $W$). Recall from the formula 7 that
$Y-m^{*}(X,Z,W)=(W-e_{W}(X))X\gamma+\varepsilon-\mathbb{E}[\varepsilon\text{ }|\text{ }X,Z,W].$
Under heteroscedasticity, we can additionally use the pilot study to estimate $\sigma^{2}(X)$. This can be done simply by noting that
$Y-m(X,Z,W)=\varepsilon-\mathbb{E}[\varepsilon\text{ }|\text{ }X,Z,W],$
where $m(X,Z,W)=\mathbb{E}[Y\text{ }|\text{ }X,Z,W]$, and so we need only estimate the variance of $Y-m(X,Z,W)$ as a function of $X$. Letting $\hat{\sigma}^{2}(X)$ be such a function and $\hat{\Sigma}$ be the diagonal matrix whose entries are $\hat{\sigma}^{2}(X_{i})$, this immediately suggests the weighted least squares estimator:
$\hat{\gamma}_{WLS}=(X^{\top}\hat{D}\hat{\Sigma}^{-1}\hat{D}X)^{-1}X^{\top}\hat{D}\hat{\Sigma}^{-1}Y.$
Likewise, the corresponding objective function to minimize is simply
$\hat{V}_{\mathrm{approx}}(e_{Z})=\bar{X}_{n}^{\top}\left(X^{\top}\text{diag}\left(\frac{\hat{e}_{W}(X)(1-\hat{e}_{W}(X))}{\hat{\sigma}^{2}(X)}\right)X\right)^{-1}\bar{X}_{n}.$
If we allow $\sigma^{2}(X,W)$ to vary with both $X$ and $W$, a non-convex objective can result in some cases. However the following theorem provides a condition under which convexity is preserved.
Theorem 3.
If $\hat{\sigma}^{2}(X,W)$ varies with both $X$ and $W$ and the condition
$\frac{1}{2}\leq\frac{\hat{\sigma}^{2}(X_{i},1)}{\hat{\sigma}^{2}(X_{i},0)}\leq 2$
holds for all $X_{i}$ with $1\leq i\leq n$, then the weighted least squares procedure yields a convex objective function $\hat{V}_{\mathrm{approx}}(e_{Z})$.
Proof.
We focus our attention to the objective $V_{\mathrm{approx}}(e_{Z})$ rather than $\hat{V}_{\mathrm{approx}}(e_{Z})$, since the latter merely substitutes in estimated functions for oracle ones. The final conclusion will then hold with $\sigma^{2}(X,W)$ replaced by $\hat{\sigma}^{2}(X,W)$. In addition, $V_{\mathrm{approx}}$ is convex in $e_{Z}(X)$ if and only if it is convex in $e_{W}(X)$ since the latter is a linear function of the former with positive coefficients, and so we can focus on convexity in $e_{W}(X)$.
To see the possible non-convexity, recall from the derivation in Section 3.3 that we must move a conditional expectation inside of a matrix inverse when switching from $V(e_{Z})$ to $V_{\mathrm{approx}}(e_{Z})$, and that this expectation passes to the innermost diagonal matrix. In the fully heteroscedastic setting, a term in this final expectation is
$\begin{split}g_{i}(e_{W}(X_{i}))&:=\mathbb{E}\left[\frac{(W_{i}-e_{W}(X_{i}))^{2}}{\sigma^{2}(X_{i},W_{i})}\text{ }|\text{ }X_{i}\right]\\ &=e_{W}(X_{i})\frac{(1-e_{W}(X_{i}))^{2}}{\sigma^{2}(X_{i},1)}+(1-e_{W}(X_{i}))\frac{e_{W}(X_{i})^{2}}{\sigma^{2}(X_{i},0)}.\end{split}$ (16)
The variance function $V_{\mathrm{approx}}(e_{Z})$ is then
$V_{\mathrm{approx}}(e_{Z})$ $=\bar{X}_{n}^{\top}(X^{\top}\text{diag}(g(e_{W}(X))X)^{-1}\bar{X}_{n}$
$=\bar{X}_{n}^{\top}\left(\sum_{k=1}^{n}g_{i}(e_{W}(X_{i}))X_{i}X_{i}^{\top}\right)^{-1}\bar{X}_{n}.$
By the vector composition rules for convex functions (e.g., Chapter 3 of Boyd and Vandenberghe (2004)), $V_{\mathrm{approx}}$ is convex in $e_{W}(X)$ if each $g_{i}$ function as defined in 16 is concave. This cannot be guaranteed for general $\sigma^{2}(X_{i},1)$ and $\sigma^{2}(X_{i},0)$. However, there is still a wide range of allowable values of $\{\sigma^{2}(X,1),\sigma^{2}(X,0)\}$. Taking the second derivative of $g_{i}(e_{W}(X_{i}))$ with respect to $e_{W}(X_{i})$, we arrive at
$\frac{1}{\sigma^{2}(X_{i},0)}\left(6e_{W}(X_{i})-4\right)+\frac{1}{\sigma^{2}(X_{i},1)}\left(2-6e_{W}(X_{i})\right).$ (17)
We seek conditions on the variance ratio under which this is guaranteed to be non-positive for any $e_{W}(X_{i})\in[0,1]$. This is a linear function in $e_{W}(X_{i})$, so it suffices to check non-positivity at the endpoints $e_{Z}=0$ and $e_{Z}=1$. Even more extreme, we can simply check non-positivity at the endpoints $e_{W}=0$ and $e_{W}=1$:
$e_{W}(X_{i})=0:\frac{-4}{\sigma^{2}(X_{i},0)}+\frac{2}{\sigma^{2}(X_{i},1)}\leq 0\iff\frac{\sigma^{2}(X_{i},1)}{\sigma^{2}(X_{i},0)}\geq\frac{1}{2},$
$e_{W}(X_{i})=1:\frac{2}{\sigma^{2}(X_{i},0)}-\frac{4}{\sigma^{2}(X_{i},1)}\leq 0\iff\frac{\sigma^{2}(X_{i},1)}{\sigma^{2}(X_{i},0)}\leq 2.$
Hence, an allowable interval for the variance ratio is $[1/2,2]$. Moving from $V_{\mathrm{approx}}(e_{Z})$ to $\hat{V}_{\mathrm{approx}}(e_{Z})$, the conclusions simply change to ones about $\hat{\sigma}^{2}(X_{i},W_{i})$, which completes the proof. ∎
Note that the interval in Theorem 3 can be improved by instead checking the endpoints $e_{Z}(X_{i})=0$ and $e_{Z}(X_{i})=1$, which correspond to $e_{W}(X_{i})=p_{\mathrm{AT}}(X_{i})$ and $e_{W}(X_{i})=p_{\mathrm{C}}(X_{i})$, respectively. These choices result in an interval for the variance ratio that is wider but that depends in a piecewise manner on the estimates compliance probabilities. Hence, we stick with the simpler $[1/2,2]$ here.
In summary, if there is reason to believe that $\sigma^{2}(X,1)$ and $\sigma^{2}(X,0)$ differ for some values of $X$, our method can accommodate estimated variances whose ratio is not too extreme. Any deviations estimated to go beyond $[1/2,2]$ can be regularized to lie in this interval, which may still be useful in the design stage. To suggest such a regularization, we first assume that $\min(\sigma^{2}(X,1),\sigma^{2}(X,0))>0$. Then we will set the smaller variance to some value $\tau^{2}=\tau^{2}(X)$ and the larger one to $2\tau^{2}$. Because estimates are generally weighted in inverse proportion to variance, we can preserve the total importance $1/\sigma^{2}(X,0)+1/\sigma^{2}(X,1)$ of some value $X$ by solving
$\frac{1}{\tau^{2}}+\frac{1}{2\tau^{2}}=\frac{1}{\sigma^{2}(X,0)}+\frac{1}{\sigma^{2}(X,1)},$
which yields
$\tau^{2}=\frac{3/2}{1/\sigma^{2}(X,0)+1/\sigma^{2}(X,1)}.$
Appendix C Simulation details
All covariates were standardized to have mean zero and variance one at the outset to avoid a coefficient vector $\gamma$ with one or two substantially larger entries. In addition, because the risk score was so heavy-tailed, we opted to convert to one that is evenly spaced in $\{\frac{1}{n+1},\dots,\frac{n}{n+1}\}$ by taking the ordinal rank of a unit’s risk score. We held out $20\%$ of the data in each simulation to be used in the pilot study.
Given data $X$ and nudge propensity $e_{Z}(X)$, our data-generating procedure is as follows:
$Z\text{ }|\text{ }X$ $\sim e_{Z}(X),$
$\mathcal{C}\text{ }|\text{ }X$ $\sim\text{Cat}(\{p_{\mathrm{NT}}(X),p_{\mathrm{C}}(X),p_{\mathrm{AT}}(X)\}),$
$W\text{ }|\text{ }\mathcal{C},Z$ $=Z\mathds{1}\{\mathcal{C}=\mathrm{C}\}+\mathds{1}\{\mathcal{C}=\mathrm{AT}\},$
$Y\text{ }|\text{ }X,W,C$ $\sim f(X)+WX\gamma+10\times\mathds{1}\{\mathcal{C}=\mathrm{AT}\}-10\times\mathds{1}\{\mathcal{C}=\mathrm{NT}\}+N(0,20).$
Here, $\mathcal{C}$ is a unit’s compliance status (either complier, always-taker, or never-taker). In addition, $f(X)$ and $\gamma$ are fit from a held-out subset of the data as described in Section 4. $f(X)$ is a random forest model, while $\gamma\in\mathbb{R}^{7}$ has the following values:
[…]
With this $\gamma$, the true final LATE is $\tau_{\mathrm{LATE}}=10.6$. The error term added to $Y$ depends on the unknown compliance status and is designed such that always-takers receive a boost in their length of stay and never-takers receive a drop. Letting $r(X)$ be a patient’s risk score, we chose the compliance probabilities as
$p_{\mathrm{AT}}(X)$ $=0.8e^{-5r(X)}+0.05$
$p_{\mathrm{AT}}(X)$ $=0.8e^{-5(1-r(X))}+0.05$
$p_{\mathrm{C}}(X)$ $=1-p_{\mathrm{AT}}(X)-p_{\mathrm{NT}}(X).$
These are visualized in Figure 5. They incorporate an intuitively reasonable expectation that higher risk patients are much more likely to be always-takers and lower risk patients are much more likely to be never-takers.
In addition, Figure 6 compares the values of the objective function at each optimal constrained solution to provide a sense of how much each constraint set inflates the optimal (approximate) variance. The unconstrained (base) optimum gives an objective value of roughly $9.47\times 10^{-4}$. For ease of interpretability, Figure 6 plots the ratio of a given solution’s objective value to this quantity. The monotonicity constraint only results in about a $7\%$ inflation of the objective, whereas using all three constraints with $\rho=0.95$ results in about an $83\%$ inflation.
Figure 5: Compliance probabilities as a function of transformed risk score. In this simulation, patients with very low risk are more likely to be never-takers, those with very high risk are more likely to be always-takers, and those in the middle are more likely to be compliers.
Figure 6: Value of the objective function under each set of constraints, normalized by the value of the objective function at the optimal unconstrained solution.